OpenStudy (anonymous):

The sum of two numbers is 50 and the difference is 18 . What are the numbers?

OpenStudy (anonymous):

@NietzscheAlex

OpenStudy (g_osmenaj):

For this one, you have to play around with different numbers until you find 2 numbers that add up to 50 and when you substract them, you get a difference of 18.

OpenStudy (phi):

is this algebra? give the unknown numbers names: people use x and y a lot the sum of x and y is 50. write that as x+y= 50 the difference is 18. Assume that x is the bigger number. then x-y= 18

OpenStudy (phi):

you have 2 equations and 2 unknowns x+y=50 x-y=18 I would add the two equations together. that means add the left sides and add the right sides. put an = sign in between

OpenStudy (anonymous):

@phi oh okay let me try

OpenStudy (anonymous):

@NietzscheAlex

OpenStudy (anonymous):

@puppylove<3123

OpenStudy (anonymous):

@janaixo